# Weak lensing of strong lensing: beyond the tidal regime

### v2

@@ 2.1 Partial displacement, convergence, shear and flexion
where ^$\hat{\mathcal{F}}_{l}$ and $\hat{\mathcal{G}}_{l}$ are the two flexion types. The effect of these quantities on an Einstein ring produced by a simple lens model consisting of a Singular Isothermal Sphere (SIS, [22]) is shown in fig. 2.
@@
Here there are only two independent complex flexion parameters, because the other third derivatives of the potential are just the conjugates of the above definitions. We then define the partial flexion parameters, which can be linked to the ones defined above, as
@@ 3.2 Source-position transformation
Let us now highlight several points. First, we can see that the two types of LOS flexion defined here ($\mathcal{F}_{\rm LOS}$ and $\mathbb{F}_{\rm LOS}$) have very different roles. The first one, just like the LOS shear, is mathematically equivalent to an external perturbation to the main lens within the same plane. On the other hand, the second type of LOS flexion cannot be reabsorbed in the main lens plane; it is inherent to the line of sight. This is very different from the tidal approximation developed in [15], where one stops at the level of the shear. In the tidal approximation, starting from a main lens and a set of perturbers along the line of sight, we are able to bring all the perturbations into the main lens plane. This is impossible in the flexion regime: we will always be left with some LOS flexion. This means that $\mathbb{F}_{\rm LOS}$ and $\mathbb{G}_{\rm LOS}$ are not degenerate with the main lens or any substructure of the main lens. An illustration of the effect of the subsequent transformations is given in fig. 3.
@@ 4 Measuring LOS flexion in strong lensing images
At this stage, a natural question to ask is whether the LOS flexion can be measured from strong-lensing images, as it could give us extra information about the matter distribution around the line of sight, especially since flexion is sensitive to the small scale distribution of matter [18]. Of course, this will depend on its magnitude; if the LOS flexion is too small, it will have a correspondingly small effect on a given lensing image, and we might not be able to measure it. To assess the measurability of the LOS flexion, we generate strong lensing images using the software lenstronomy [28, 29], limiting our analysis to Einstein rings, and perform Markov chain Monte Carlo (MCMC) parameter inference to recover the parameters that were used to create the image in the first place, similarly to what was done in [16]. To do this, we have implemented the LOS flexion formalism in lenstronomy. The code is currently available on Github in a branch called los-flexion. [^2] https://github.com/TheoDuboscq/lenstronomy/tree/los-flexion. The use of the LOS flexion in lenstronomy is similar to that of the LOS shear, with the user simply needing to call the relevant lens model profile and define the corresponding LOS parameters.
@@ 4.1 The expected value of cosmic flexion
It turns out, however, that according to the above definitions the variance of flexion diverges. Since the flexion is one derivative higher than the shear, it strongly depends on small scales. This results in an additional $\ell^{2}$ in the expression of the variance. [^3] The shear variance has a similar form to the flexion variance, but with a different kernel $q(\chi)$ and $\ell$ instead of $\ell^{3}$. This means that flexion is highly sensitive to high $\ell$, and if we extrapolate $P_{\delta}$ to high $\ell$ using a power law, the variance diverges. Thus, the value we obtain strongly depends on the upper bound we set in the numerical integral.
@@
This gives us a guideline for our mock images, as it gives us the expected scale the flexion should have. As previously stated, the scale $\theta_{S}$ we average over should correspond to the scale of the lensed images. We choose this scale to be the effective Einstein radius of the main lens. Of course, when taking $\theta_{S}$ to zero the variance of the flexion diverges, as it is equivalent to not taking the average anymore, hence recovering eq. 4.6. Looking at those curves, we see that the variance of the $os$ flexion is greater than the others. This is due to the kernel $W_{os}(\chi,\bm{\Pi})$ involved in the expression of the variance (fig. 12 in appendix A shows a comparison of the different flexion kernels). Because of this, one might be tempted to only keep the $os$ flexion terms and neglect the others. We should refrain from doing so; first because there is only a factor 3 between their standard deviation, second because this would mean taking $\mathbb{F}_{\rm LOS}$ and $\mathbb{G}_{\rm LOS}$ to 0, losing a key term of the lens equation. We therefore do not make such an approximation, and keep all the terms for the subsequent analysis.
@@ B.2 Averaged flexion
Notice that unlike the integral in eq. B.8, this one converges. Indeed, $P_{\rm F}(\ell)\in\Theta(l^{-2})$ and $J_{1}(\ell\theta_{S})/\ell\in\Theta(l^{-3/2})$ for $\ell\to\infty$, so the integrand is of order $\ell^{-2}$ which means the integral converges.

### v3

@@ 2.1 Partial displacement, convergence, shear and flexion
where ^$\hat{\mathcal{F}}_{l}$ and $\hat{\mathcal{G}}_{l}$ are the two flexion types. Here there are only two independent complex flexion parameters, because the other third derivatives of the potential are just the conjugates of the above definitions. The effect of these quantities on an Einstein ring produced by a simple lens model consisting of a Singular Isothermal Sphere (SIS, [22]) is shown in fig. 2.
@@
To generate those images, we used the software lenstronomy [23, 24], in conjunction with the LOS flexion formalism developed in section sec. 3, where we define the notion of LOS convergence, shear and flexion. For each panel, we put a non zero magnitude for one of these LOS parameters. An implementation of this model in lenstronomy is available on Github in a branch called los-flexion. [^2] https://github.com/TheoDuboscq/lenstronomy/tree/los-flexion. We then define the partial flexion parameters, which can be linked to the ones defined above, as
@@ 3.2 Source-position transformation
Let us now highlight several points. First, we can see that the two types of LOS flexion defined here ($\mathcal{F}_{\rm LOS}$ and $\mathbb{F}_{\rm LOS}$) have very different roles. The first one, just like the LOS shear, is mathematically equivalent to an external perturbation to the main lens within the same plane. On the other hand and quite remarkably, the second type of LOS flexion cannot be reabsorbed in the main lens plane; it is inherent to the line of sight. This is very different from the tidal approximation developed in [15], where one stops at the level of the shear. In the tidal approximation, starting from a main lens and a set of perturbers along the line of sight, we are able to bring all the perturbations into the main lens plane. This is impossible in the flexion regime: we will always be left with some LOS flexion. This means that $\mathbb{F}_{\rm LOS}$ and $\mathbb{G}_{\rm LOS}$ are not degenerate with the main lens or any substructure of the main lens, suggesting a way to distinguish LOS perturbers from haloes within the main lens plane. This possibility will be further investigated in future work. An illustration of the effect of the subsequent transformations is given in fig. 3.
@@ 4 Measuring LOS flexion in strong lensing images
At this stage, a natural question to ask is whether the LOS flexion can be measured from strong-lensing images, as it could give us extra information about the matter distribution around the line of sight, especially since flexion is sensitive to the small scale distribution of matter [18]. Of course, this will depend on its magnitude; if the LOS flexion is too small, it will have a correspondingly small effect on a given lensing image, and we might not be able to measure it. To assess the measurability of the LOS flexion, we generate strong lensing images using lenstronomy, limiting our analysis to Einstein rings, and perform Markov chain Monte Carlo (MCMC) parameter inference to recover the parameters that were used to create the image in the first place, similarly to what was done in [16]. To do this, we use the LOS flexion formalism we have implemented in lenstronomy mentioned in sec. 2. The use of the LOS flexion in lenstronomy is similar to that of the LOS shear, with the user simply needing to call the relevant lens model profile and define the corresponding LOS parameters.
@@ 4.1 The expected value of cosmic flexion
Our results are consistent with the ones first obtained in [18, 30] in the context of weak lensing. As a matter of fact, the flexion formalism was naturally first introduced in the context of cosmic weak lensing in [18]. The only difference lies in the specific shape of the weight function $W$; the weak lensing scenario corresponds to $W_{os}$ in the present paper.  [^3] Beware, our definition of $\mathcal{F}$ and $\mathcal{G}$ differs from the ones of [18, 30] by a factor $1/2$, which explains the factor 1/4 in our flexion power spectrum. It turns out, however, that according to the above definitions the variance of flexion diverges. Since the flexion is one derivative higher than the shear, it strongly depends on small scales. This results in an additional $\ell^{2}$ in the expression of the variance. [^4] The shear variance has a similar form to the flexion variance, but with a different kernel $q(\chi)$ and $\ell$ instead of $\ell^{3}$. This means that flexion is highly sensitive to high $\ell$, and if we extrapolate $P_{\delta}$ to high $\ell$ using a power law, the variance diverges. Thus, the value we obtain strongly depends on the upper bound we set in the numerical integral.
@@
This gives us a guideline for our mock images, as it gives us the expected scale the flexion should have. As previously stated, the scale $\theta_{S}$ we average over should correspond to the scale of the lensed images. We choose this scale to be the effective Einstein radius of the main lens. Of course, when taking $\theta_{S}$ to zero the variance of the flexion diverges, as it is equivalent to not taking the average anymore, hence recovering eq. 4.6. Looking at those curves, we see that the variance of the $os$ flexion is greater than the others. This is due to the kernel $W_{os}(\chi,\bm{\Pi})$ involved in the expression of the variance (fig. 12 in appendix A shows a comparison of the different flexion kernels). Because of this, one might be tempted to only keep the $os$ flexion terms and neglect the others. We should refrain from doing so; first because there is only a factor 3 between their standard deviation, second because this would mean taking $\mathbb{F}_{\rm LOS}$ and $\mathbb{G}_{\rm LOS}$ to 0, losing a key term of the lens equation. We therefore do not make such an approximation, and keep all the terms for the subsequent analysis. Finally, let us point out that this method is not the only renormalization solution. The authors of [30] propose, based on [31], an alternative approach. They compute the power spectrum up to a large $k_{max}$ and then impose a steeper slope of the power spectrum above this threshold, resulting in the convergence of the integral. With this method, they find a variance quite similar to ours; looking at fig. 3 of [30] we find $\sigma_{\mathcal{F}}=\sigma_{\mathcal{G}}\approx 7\times 10^{-3}\,\rm arcsec^{-1}$, which is close to our $\sigma_{\mathcal{F}_{os}}=\sigma_{\mathcal{G}_{os}}$ for the $\theta_{S}$ of a typical strong lensing image ($\theta_{S}\sim 1$ arcsec).
@@ B.2 Averaged flexion
Note that unlike the integral in eq. B.8, this one converges. Indeed, for high $\ell$ we extrapolate $P_{\rm F}(\ell)\in\Theta(l^{-3})$ and $J_{1}(\ell\theta_{S})/\ell\in\Theta(l^{-3/2})$ for $\ell\to\infty$, so the integrand is of order $\ell^{-3}$ which means the integral converges.